OpenStudy (anonymous):

Need Help with 10 questions- Evaluating Functions.

OpenStudy (anonymous):

Question 1 If f(x) = x + 2, find f(3). 1 6 5 −1 -------------------------------------------------------------------------------- If f(x) = −4x − 1, find f(−2). −9 −7 −10 7 -------------------------------------------------------------------------------- Question 3 Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie. c(x) = 2.00x + 0.50 c(x) = 2.00 + 0.50x c(x) = 2.50x c(x) = (2.00 + 0.50)x -------------------------------------------------------------------------------- Question 4 Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura keeps the movie for 3 nights, how much will she pay when she returns it? $2.50 $6.50 $ 3.50 $7.50 -------------------------------------------------------------------------------- Question 5 If g(x) = x2 + 2, find g(3). 9 8 11 6 -------------------------------------------------------------------------------- Question 6 Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura returned the movie and paid $5.00, how many nights did she keep the movie? 14 1.5 8 6 -------------------------------------------------------------------------------- Question 7 If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10 -------------------------------------------------------------------------------- Question 8 If h(x) = 2x + 1, find the value of x if h(x) = 11. 5 23 6 10

OpenStudy (anonymous):

Question 1: Substitute in your x, which is 3. f(3)=3+2 f(3)=5 Question 2: Substitute in your x, which is -2 f(-2)= −4(-2) − 1 f(-2)=8-1 f(-2)=7

OpenStudy (anonymous):

Question 5: g(x) = x^2 + 2, find g(3). g(3)=(3)^2+2 g(3)=9+2 g(3)=11 Question 7: If g(x) = 2(x − 4), find the value of x if g(x) = 20. 20=2(x-4) 20=2x-8 28=2x x=14

OpenStudy (anonymous):

Thank you so much could you help me with the other 6 questions please??

OpenStudy (anonymous):

Question 8: If h(x) = 2x + 1, find the value of x if h(x) = 11 11=2x+1 10=2x x=5

OpenStudy (anonymous):

Question 3: Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie. c(x) = 2.00 + 0.50x

OpenStudy (anonymous):

just need question 4 and six

OpenStudy (anonymous):

Question 4 Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura keeps the movie for 3 nights, how much will she pay when she returns it? You use the equation that was derived from Question 3. c(x)=2.00+0.50x c(x)=2.00+0.50(3) = $ 3.50

OpenStudy (anonymous):

Try them on your own and tell me what you get, if your stuck I'll help you.

OpenStudy (anonymous):

correct ans help